Find the derivative of the inverse of f(x) by using 1/derivative of f(x).
1. F(x)=(x+3)^.5
This problem does not sound correct to me, but that is the problem, please help.
Thanks
: Find the derivative of the inverse of f(x) by using 1/derivative of f(x).
1. F(x)=(x+3)^.5
This problem does not sound correct to me, but that is the problem, please help.
Thanks

Answer by algebrapro18(155)

Find the derivative of the inverse of f(x) by using 1/derivative of f(x).
1. F(x)=(x+3)^.5
This isn't right.
lets first find the inverse of f(x) and then find the derivative of it.
y = (x+3)^(1/2)
x = (y+3)^(1/2)
x^2 = y+3
x^2 - 3 = y
so y'= 2x
now lets find the derivative of x and then take 1/y'.
y = (x+3)^(1/2)
y' = (1/2) * (x+3)^(-1/2) * 1 (by the chain rule)
y' = 1/(2(x+3)^(1/2))
1/y' = 2(x+3)^(1/2)
but as you can see 2x does not equal 2(x+3)^(1/2)